Scientific Justification

Introduction The accelerating expansion of the universe is one of the more captivating “key scientific questions of our day” identified by the NRC Committee on the Physics of the Universe (Tur02). In the few years since the acceleration was first seen in the Type Ia supernova (SN Ia) Hubble diagram (Per99, Rei98), the evidence has grown even stronger: complementary CMB measurements have indicated that the Universe has zero curvature (Sie03), making the SN Ia result more determinative, and – in combination with the SNe — pointing to a [Ω_{M} ≈ 0.3, Ω_{Λ} ≈ 0.7] cosmology. This is also consistent with other astronomical mass density measurements (Perc02).

The SN Hubble diagram remains the only direct approach currently in use to study acceleration. There are ongoing studies of the details of all known relevant sources of systematic uncertainty, but none show any biases at a level that might affect the basic acceleration results. These include: any changes with z in host-galaxy extinction by ordinary dust (Per99,Rei98,Sul03), any extinction by intergalactic gray dust (Pra03), gravitational-lensing (de)amplification of SN magnitudes (Per99), discovery selection effects (Per99,Rei98), K-correction systematics (Nug02), and population drifts in the SN environment (Sul03).

With these advances in hand, it is now time to pursue the cause of the acceleration — the “dark energy” — be it a simple “energy of the vacuum” (Einstein’s cosmological constant, Λ) or a general dynamical scalar field (as is assumed responsible for inflation). Although extremely challenging, we can in fact address this scientific question in several, ordered steps: large projects are just now starting which will test the hypothesis that the expansion is consistent with Λ, which is characterized by a constant equation of state ratio w ≡ p/ρ = −1. Later projects will tackle the still more difficult goal of detecting changes in w indicative of scalar field models. (The satellite experiment, SNAP has been proposed for this work.)

The separate proposal to study SNe Ia at z ∼ 0.5 submitted by our group for Cycle 12 is aimed at the first key step, testing Λ. However, since there is a Cycle 12 Treasury proposal from Faber et al. that includes a wide-area Groth-field search for SNe at 1 < z < 1.7, we are also submitting this non-proprietary proposal since there are not many opportunities available to do the exploratory work at these higher redshifts that will lay the necessary groundwork for future steps in the study of dark energy. We are especially interested in exploiting the synergy of such an HST SN search with our work on the same region as part of the SNLS. If the “Tier 3” SN search of the Faber et al. proposal is approved, that effort and the follow-up of the resulting supernovae represent a total level of investment of HST resources well beyond what we would propose for a regular GO proposal, especially as an addition to our z ∼ 0.5 proposal. We have therefore designed and will now describe a non-proprietary program in the spirit of a Treasury program to economically follow up each of the supernovae: we would propose to organize this follow-up and obtain the data for the astronomical community. All the data would be available to everybody immediately, and we would welcome any others who would want to work with us on this data. The goal would be to treat this follow-up work as a service to the community, and deliver the “standard-ized candle” data that will be important for the community to obtain for these supernovae. If it would be considered more appropriate, we would be happy to run this follow-up as part of a Treasury or Director’s Discretionary program.

We have discussed this proposal with Riess, et al., who are also proposing to follow these same supernovae, and both teams recognize the advantages of comparing several approaches to the SN selection and follow-up strategy. The two teams are therefore willing to each observe a fraction of the SNe Ia discovered in Cycle 12, yielding the “right of first refusal” to the other team based on a prearrangement of preferred fields (or any other system the TAC deems appropriate). This will allow a reduction in the telescope allocation needed to achieve fully independent results from the two teams’ approaches, as has been so valuable previously. If we share in the allocation for following SNe Ia found by HST, we would also pool our efforts in searching for the supernovae, so that every supernova that could be found will be found (and followed), taking advantage of techniques of both groups. Simple Scientific Goals If ∼10 new SNe Ia were added to the sample at 1 < z < 1.7, it would almost triple the number of HST-discovered SNe available for studies in this redshift range. In all, it would become possible to populate the Hubble diagram in this redshift range with over 20 SNe Ia, after adding ∼ 5 more ground-based discoveries at 1 < z < 1.3 that we have followed up with HST. This sample size would allow a more rigorous study of the universe’s expansion during a time when it is expected to be decelerating, since the mass density should be dominant over the dark energy. While a half dozen or so SNe provide a suggestive result, a score of SNe makes it possible to begin to look for trends and outliers. (In fact this first, real Hubble diagram extending well into the “epoch of deceleration” is likely to be included in the textbooks as part of the most complete record to date of the history of the expansion of the universe.)

At redshifts significantly greater than z = 1 the hunt for trends and outliers is more than an academic exploration. Although the question of gray dust has been well answered by recent Chandra measurements (Pae03), evolution in the properties of SNe Ia is still a much debated question. If z ∼ 0.5 SNe Ia are fainter due to evolution rather than an accelerating expansion, they should continue to get fainter at even higher redshifts. Since the dark energy cosmology predicts that SNe Ia will be brighter than for an open Λ = 0 cosmology, it is possible to test the hypothesis of open, Λ = 0 with evolution versus that of dark energy. With dust being less of a concern, this test of evolution is more powerful than ever because the possibility of a dust/evolution conspiracy has become remote. In addition to this evolution test, we expect to see an asymmetric distribution of residuals from the Hubble line due to the frequent de-amplification, and rarer amplification, of the supernovae’s brightnesses by the gravitational lensing of intervening mass density fluctuations (especially amplification by galaxy clusters or deamplification by their absence). This residual distribution will be important to compare with the evidence for overdense and underdense lines of sight to the supernovae — witness the many papers which have tried to understand the lensing of SN 1997ff (Mor01,Ben02). In this way, we can begin to test models of the mass distribution and power spectrum in a unique way.

The SN Ia evidence for the acceleration of the universe’s expansion can be characterized as resting on the observation that a set of SNe Ia around z ∼ 0.5 appear to be fainter than would be expected in an empty universe, a universe with no mass and no cosmological constant. It is therefore of obvious interest to see the epoch even further back in time when the SNe should appear to be brighter than expected in an empty universe. The proposed dataset, after correcting for the statistical effects of the gravitational lensing (de)amplifications will provide a much more solid basis for this simple “sanity check” for our cosmological model. Setting the Stage for Future Studies The proposed data set of supernova magnitudes should also be understood as a pilot sample to open up many years of future studies. Supernovae at these redshifts are extremely difficult to study even with the current HST instrumentation, so we do not expect to obtain the final, definitive measurements for these SNe, but rather to provide important hints as to how future studies with next-generation instruments should be designed. For example, the WFC3 will be an obvious tool for the SN work in this redshift, with a much wider-field infrared detector. There will not be time for very many major surveys to be performed with this instrument – necessary starting points for any SN studies – so it will be extremely important to make every large survey count. At the most basic level, the current proposed dataset will help establish SN rates in this redshift range, not only for the SNe Ia that are being followed up, but also for the comparable numbers of core-collapse SNe II and Ib/c that will be screened out from the follow up work. (Of course, this will also help current star-formation-history studies.) We will certainly test techniques to screen the non-Ia events that will need to be used in future studies.

In the longer run, we need to test and further develop approaches to constraining systematic uncertainties for supernovae at these extremely high redshifts. This is necessary if we are to reach the more demanding precision requirements of the future studies of changes in the dark-energy equation of state. While this work will ultimately require more advanced instrumentation (such as proposed for SNAP), the ground-work needs to be laid by establishing the simple properties of the supernovae in this redshift range, including magnitudes, colors, and timescales. It is important to note that although the dark energy is not a dominant component at these redshifts, the measurement of the Hubble diagram will need to extend through exactly this range to achieve the best sensitivity to the time variation of the equation of state. This is shown in Figure 1. If the current study-phase schedules hold, there are not many more years remaining before the design of the SNAP satellite and mission will be frozen, so this is a particularly timely study. The Desired Dataset (and its Potential Discontents) We wish to obtain the dataset of follow-up photometry for each SN Ia discovered that is likely to be of most use to the astronomical community, and do so in the most efficient way possible. The obvious prime interests are (1) to spend orbits primarily on the Type Ia SNe, not on screening out the other types, and (2) to obtain a “standardized candle” magnitude that is corrected for any extinction and for the lightcurve width-brightness relation. To accomplish this, there are a number of important constraints that must be taken into account:

1. In the Faber et al. proposed Tier 3 survey, the candidate supernovae will appear as bright points on the subtraction of one orbit of I band data from another 45 days later. These must be confirmed by at least one more image, because the high HST cosmic-ray rate will cause some coincidence false alarms in such a large survey, even with multiple cosmic-ray splits in the discovery orbit.

2. Of the confirmed supernovae, over 60% will be non-Type Ia (core-collapse) SNe. In a few cases these can be screened out using the Faber et al. proposed V band observation, since Type Ia at peak brightness show no restframe UV flux, while Type II at peak brightness are strong in the restframe UV. However, it is important to note that after 5 restframe days past peak, the Type II SNe are indistinguishable in the restframe UV from the Type Ia (see Figure 2), so a different screening test must be used – and most of the SNe II in this type of search will be discovered beyond this lightcurve date. We therefore propose to use a combination of host galaxy color constraints (from ground-based data) to identify SN host galaxies in the appropriate redshift range without high star-formation rates (see Description of Observations).

3. To establish a “standardizable candle” magnitude, it is necessary to establish the date of maximum (or some other fiducial date) on the supernova lightcurve. This will require at least two more dates of SN observation beyond the Faber et al. proposal’s discovery data points, which only sample a single date on the lightcurve. By adding one to two more points it will be possible to measure the timescale of the lightcurve (parameterized either by the stretch factor, s, or Δm_{15}), and then “correct” the supernova magnitude using the lightcurve width-brightness relation.

4. At redshifts above z = 1, the host galaxy light is usually closely associated with the supernova light, so they cannot be reliably disentangled without subtracting off an image observed a year later, after the SN has faded. This final galaxy image requires an exposure length matched to the longest exposure taken of the supernova. Note that even with this exposure length, all signal-to-noise calculations for the supernovae will be reduced, to account for the extra noise introduced in subtracting the galaxy image. (This degradation is not quite a factor of \sqrt{2}, because the noise from the SN signal itself does not enter in twice.) We have included these final galaxy images in the orbit count requested in this proposal, although they will not actually be scheduled in this Cycle, but a year later.

5. The SNe that are not found in E/S0 host galaxies will require color measurements to correct for extinction. Since the color measurement is a subtraction of two magnitudes, each of which in turn is a subtraction of two images (one with and one without the SNe), the color measurement signal-to-noise is a factor of 2 below that of an isolated star with the same magnitude as the SN. Moreover, this large uncertainty gets increased by the R_{B} ratio of color to extinction. The color measurements therefore become the ones that require the most signal to noise, i.e. the longest exposures both for lightcurve maximum and for the host galaxy image.

6. The SNe discovered in E/S0 host galaxies are particularly important for this work. First, they are very unlikely to be anything but a Type Ia (so far, only SNe Ia have been seen in ellipticals). Second, they are unlikely to have significant extinction. Since the major source of “standardized candle” magnitude uncertainty for the SNe Ia in spiral galaxies will be the extinction correction, the SNe in E/S0’s will provide the best points on the Hubble diagram, as we have demonstrated in Sul03. We take best advantage of this chance to obtain much higher precision measurements by using longer exposures for points on the elliptical-hosted SN lightcurves. This extra investment shows up quite directly as improved uncertainty on the Hubble diagram; a comparable improvement for the spiral-hosted SNe would be prohibitively expensive. The expected uncertainties will be as follows:

Follow-up Cases and Orbit Counts

z Host Type # of Orbits # of Orbits Orbit Measurement
SN+gal F110W SN+gal F160W Totals Uncertainty
1.20 spiral, etc  5+2  2+2  11  0.27 mag
1.70 spiral, etc  12+4  3+3  21  0.29 mag
1.20  E/S0  5+2  0  7  0.11 mag
1.70  E/S0  9+4  0  13  0.18 mag

Note that these uncertainty estimates do not include any dispersion that is intrinsic to SNe Ia after the lightcurve width and extinction corrections are applied. If these intrinsic dispersions are included, effective weight for a cosmology measurement for these SNe would correspond to an uncertainty between 0.15 and 0.21 mag for the E/S0-hosted SNe at redshifts between 1.2 and 1.7; the uncertainties for the spiral galaxy-hosted SNe at these same redshift extremes would change very little.

With these six constraints in mind, we have constructed an optimized follow-up observing program that targets the ∼10 SNe Ia expected from the Faber et al. Treasury proposal (Tier 3). The program assumes a spread of redshifts between 1.2 and 1.7; this is the redshift range that is least-well studied using ground-based supernova discoveries. An approximate fraction of SNe Ia in E/S0’s is taken to be 30%, based on lower-redshift experience. The actual choice of number of orbits for a given supernova will be made depending on the estimated host-galaxy redshift and the elliptical/spiral classification (both from ground-based host color data).

Total Treasury Type Ia Follow-up Program

z  Follow-up Case  Orbits  Cycle 12 Cycle 13 Total Uncertainty
Range  per Range Orbits  Orbits Orbits (E/S0, spiral)
1.20–1.35 1 E/SO + 3 spiral 7 + 3×11  26  14  40  0.11, 0.27
1.35–1.55 1 E/SO + 2 spiral 10 + 2×16  26  16  42  0.15, 0.28
1.55–1.70 1 E/SO + 2 spiral 13 + 2×21  37  18  55  0.18, 0.29
89  48  137 Total Orbits

References (Ben02) Benítez, N., et al. 2002, ApJ, 577, L1. (Cap99) Cappellaro, E., Evans, R., & Turatto, M. 1999, A&Ap, 351, 459. (Hut01) Huterer, D. & Turner, M., 2001, Phys. Rev. D64, 123527 (Nug02) Nugent, P., A. Kim, S. Perlmutter, 2002, PASP, 114, 803. (Mor01) Mörtsell, E. et al. 2001, ApJ, 561, 106. (Pae03) Paerels, F., et al. 2003, ApJL, submitted. (Per97) Perlmutter, S., et al. 1997, ApJ, 483, 565. (Per99) Perlmutter, S., et al. 1999, ApJ, 517, 565. (Perc02) Percival, W., et al. 2002, MNRAS, 337, 1068. (Phi99) Phillips, M., et al. 1999, AJ, 118, 1766, (Rie98) Riess, A. G., et al. 1998, AJ, 116, 1009.

Figure 1: Comparisons of simulated measurements of the time variation of w, the dark-energy equation of state ratio. The measurements are based on a simulated Hubble diagram for SNe Ia that extends out to redshift z_{max}. The uncertainty in this measurement, σ(w′) improves significantly as the redshift range of the SN sample increases, until it levels off around z = 1.7 for a realistic scenario with systematic uncertainties (at a level achievable with future instruments currently being designed), and a prior on Ω_{M} with 0.03 uncertainty. Less realistic scenarios with even better prior knowledge of Ω_{M} (and possibly an idealized supernova and idealized measurements, i.e. no systematic uncertainties) give uncertainties in w′ that level off around z = 1.5 or 2.3.

(Sie03) Sievers, J. L, et al. 2003, ApJ submitted. astro-ph/0205387 (Sul02) Sullivan, M., et al. 2000, MNRAS, 319, 549. (Sul03) Sullivan, M., et al. 2003, MNRAS, in press. astro-ph/0211444 (Tur02) Turner, M.S., et al. 2002, Report of the Committee on the Physics of the Universe, National

Academies Press

Description of the Observations

The Search: In coordination with the Faber et al. Treasury team we will plan the execution of the SN search in a way that optimizes the number of SNe Ia beyond the acceleration/deceleration transition epoch, z > 1.2, while providing discrimination against core-collapse SNe. The basic strategy is to obtain pairs of F814W images roughly 45 days apart, with a matching F606W image

Figure 2: Comparative spectral energy distributions for redshifted Type Ia and Type II SNe. The thick red line shows the maximum-light spectrum of an SN Ia redshifted to z = 1.56. The series of black curves show the spectrum of an SN II at maximum, at +5 days, and and +25 days, in this case redshifted to z = 1.50. The wavelength ranges of the Treasury search filters are overlaid. It is apparent that the F606W-F814W colors for an SN Ia at maximum and an SN II just a few days after maximum (for Type IIp) will be similar when these objects have similar redshifts. Most of the SN II will be detected after maximum (the most probable date is 10 days after maximum). Thus, Treasury F606W images — of 1 orbit duration — will generally not detect the z > 1.2 SNe II, and it will be impossible to screen out the SNe II based on the signal in this band. Fortunately, as described in the text, other methods can be used to distinguish SNe Ia and SNe II.

at the second epoch. The search gap is matched to the risetimes of z > 1.2 SNe Ia. (For example, at z = 1.5, this is an 18 day gap in the SN restframe.) The F606W image will allow only some discrimination against SNe II near maximum (see below). As the z ∼ 1.7 redshift limit of the search is reached, the range of dates around maximum light over which an SN Ia can be detected will narrow. Thus the search guarantees that the SNe Ia of interest will be near maximum light when found — a feature we can exploit in selecting candidates and scheduling follow-up. All observations will have CR-SPLITS so that the number of spurious detections is small (See below for discussion of confirming images.) As demonstrated by the 100+ SNe Ia out to z = 1.3 that we have discovered, we have extensive experience with deep ACS imaging and with SN searches using data of all types. In preparation for this project, we will tune our software with the GOODS dataset, and compare our results with the reported transient discoveries. Discrimination against core-collapse SNe: Our simulations based on current estimates of the 1 < z < 2 star formation history (Sul02) indicate that for each SN Ia, the Treasury search will detect one Type II-P and a Type II-L SN. Some Type Ic may also be detected, but their rates and spectral properties are too poorly known for a believable simulation. The UV spectra of SNe Ia are such that for z > 1.2 they will not be detected in F606W. At maximum light, bright SNe II can be detected in F606W because they are much hotter — this has been advertised as an easy way to discrimate between these two types of SNe. However, SNe IIp experience a long plateau phase, during which they are only slightly less luminous than at B-band maximum. Thus, it is more probable that the F814W detections will find an SN IIp on the plateau phase (our simulations show the most probable detection is at 10 days after B-band maximum). As shown in Fig. 2, the spectrum of a Type II will look very much like that of a Type Ia, and in particular, the Type II will not be detected in F606W. Therefore, additional techniques are required to eliminate SNe II.

The CFHT SuperNova Legacy Survey (SNLS) will be running a continuous SN search with multiband data every other night in the Extended Groth Strip during the 5-month period when the Treasury search would be scheduled. This means that most z < 1.2 SNe Ia, and z < 0.8 core collapse SNe will have several ground-based lightcurve points which can be combined with the Treasury observations to exclude the z < 1.2 SN Ia candidates for follow-up by the program proposed here. The SNLS SNe Ia in the range 1 < z < 1.2 would also be suitable as backup targets. Note that we are participants in SNLS, and thus have access to this proprietary data set for this purpose (although note that the rules of the CFHT Legacy Survey precludes this ground-based data from being made public as part of this HST proposal).

Essential additional ingredients for discriminating against core-collapse SNe are the SN redshift and the star-formation rate of the hosts. The HST/DEEP team (a collaboration closely associated with the Faber et al. Treasury proposal collaboration) will supply photometric redshifts able to discriminate z < 0.85 from z > 0.85 host galaxies. Unfortunately, such a redshift cut is insufficient for identifying the host galaxies at z > 1.2. (By the time of the Treasury search the DEEP team may also be able to supply some spectroscopic redshifts as well.) Our strategy is therefore to use deep colors from CFHT multi-band imaging of the Groth strip to provide further redshift selection and star-formation-rate selection. These colors can determine photometric redshifts for more than two-thirds of the SN host galaxies with 1.2 < z < 1.7, with improved reliability when coupled with additional deep K-band imaging being obtained by members of the DEEP team. Since core-collapse SNe occur in actively star-forming galaxies — most of which will be UV-bright and therefore detectable in the optical — we should have photometric redshifts for most of the 1.2 < z < 1.7 core-collapse SNe. In addition, CFHT colors will allow us to estimate the UV spectral slope and star-formation rates, making it possible to discriminate against actively star-forming galaxies and their attendent core-collapse SNe. We estimate that the corresponding fraction of Type Ia SNe eliminated by this same cut will be small (Cap99).

We should also note that as one of the goals is to blaze the trail for future experiments in this redshift range, knowing the rate of core-collapse SNe is important in itself. And of course, there is a great deal of interest within the astronomical community is using these SN rates to trace the evolution of the star-formation rate. HST Follow-up Strategy: The basic information needed to put a SN Ia on a Hubble diagram are a standardized, extinction-corrected magnitude and a redshift. Most methods use the lightcurve width, Δm_{15} or stretch s, to standardize SN Ia magnitudes, as follows:

M_{B_{corr}} = C + B_{max} + 1.7(s − 1) + R_{B}E(B − V)

(see Per97 or Phi99). We will use this standardization technique for all Treasury SNe, but in the case of E/S0 hosts we can assume that E(B − V) = 0. This conserves much of the exposure time needed for the restframe B-V measurements neccessary for SNe Ia in spiral hosts. Given the redshifts of these objects, the most suitable approach is to measure the lightcurve shape and peak using F110W, and then use F160W to obtain the color. (An alternative would be to obtain the lightcurve in ACS F850LP, however this filter would exclusively sample the restframe UV which is less well characterized and where our HST UV spectroscopy program has found variety among SNe Ia. Moreover, when convolved with the SN spectrum (see Fig. 2) the sensitivity in F850LP and F110W are comparable.) We would take a first epoch in F110W to obtain the very first lightcurve point, but also as a check against possible spurious candidates. We would obtain the F110W−F160W color at the second epoch, by which time the results of the initial F110W observations will have been evaluated, but the SN Ia still will be near peak brightness. With the F110W−F160W color in hand we will reevaluate candidates for their consistency with SNe Ia. If once-promising candidates are eliminated in this process, subsequent orbits in the series for that object will either be cancelled or redirected to other candidates in the field.

Special Requirements

This program is a ToO for follow-up of high-redshift supernovae to be discovered by a proposed Treasury program, PI-Faber. As such, it will be necessary to coordinate the scheduling of this proposal with the scheduling of the Treasury program. Once the timing of the Treasury program is established, we will consult with the HST schedulers concerning the likely plan for the execution of our program. We note that if the supernova search portion of the Treasury program is scheduled into a monolithic slot such that several candidate SNe are guarenteed in any given week, then our follow-up could be scheduled months in advance. All targets would be in the Extended Groth Strip, and thus within the ±3600 arcsec offsetting range of HST. This enables all details of the observations to be scheduled well in advance, and then a week before the necessary calendars are built we would provide the schedulers with the exact target location. Alternately, if the Treasury search is broken into small portions for scheduling reasons, we would try to interweave all the time-critical aspects of the follow-up observations from this program with those of the z ∼ 0.5 SN Hubble Diagram HST proposal that we have submitted, which is following SNLS SNe Ia discovered in the same field (assuming that is approved). In this mode, we would either observe a Treasury supernova or an SNLS supernova with one orbit on NICMOS, with only a change between NIC1+F110M and NIC2+F110W required. Therefore, these observations would be a very low impact ToO if the Treasury search is concentrated in time and/or if our SNLS z ∼ 0.5 SN follow-up proposal is approved.

Coordinated Observations

As described under Description of Observations, supporting photometric and/or spectroscopic redshifts of host galaxies from the DEEP team will be employed to aid to evaluating the probable redshift and type of Treasury SN candidates. Deep multi-band CFHT photometry will also be essential for selecting the high-redshift SN candidates that are Type Ia. These observations will be performed in advance of the Treasury search, and therefore should not constrain the HST schedule.

Equally important, we will have SNLS detections of many of these same SNe, including SNe Ia out to z ∼ 1.2. This will be extremely powerful in eliminating lower redshift SNe Ia and many core-collapse SNe. We would aim to schedule the Treasury search during the 5-month period during which SNLS will be observing the Extended Groth Strip in a continuous SN search mode.

Justify Duplications

All targets are unique transient objects, and therefore are not duplicated by other programs.

Previous Related HST Programs

This proposal builds on our previous HST programs that have studied ∼ 30 mid to high redshift SNe, providing multi-epoch lightcurves and images of host galaxies using WFPC2, NICMOS, STIS, and most recently, ACS. GO-9705: In this program, we pushed our SNe Ia studies to the highest redshifts that are feasable for a ground-based discovery and spectroscopic identification campaign. HST follow-up observations for this program started after servicing mission 3B in March 2002 and have been completed for the most past - final reference images are still to be taken. Coordinated with three large search campaigns using the Subaru 8.2 m and also with simultaneous smaller searches using the CTIO 4 m and CFHT 3.6 m, we obtained ACS/WFC and NICMOS/NIC2 photometry for multi-epoch lightcurves of eight Type Ia SNe at high redshift (0.9 < z < 1.3). For two of the highest redshift SNe, ACS grism spectra were taken. Analysis of this ACS data is in progress. With the refurbished NICMOS, we obtained final reference images of the host of SN1998eq, which we had previously studied in G0-8088, and these images will allow us to complete that analysis.

GO-8585: In GO 8585 we observed six Type Ia supernovae with HST using WFPC. The supernovae were discovered in ground based searches at the CTIO 4-m, CFHT and Subaru telescopes. We obtained both U- and B-band restframe photometry (using either F814W or F850LP depending on the redshift) for each supernova for a period of 2 months. Analysis of this data will be completed when the final reference images are available, scheduled for spring 2003.

GO-8313: The objective of this project, which has now been completed with a publication in press (Sul03), was to obtain snapshot unfiltered STIS images of distant galaxies of known redshift which have hosted supernovae (SNe) of Type Ia found by the SCP, 20 of which are used in the Hubble diagram of 42 type Ia SNe (Perlmutter et al. 1999). In Sullivan et al. (Sul03) we present these new results on the Hubble diagram of SNe Ia as a function of host galaxy morphology that demonstrates that host galaxy extinction is unlikely to systematically dim distant SN Ia in a manner that would produce a spurious cosmological constant. The internal extinction implied is small, even for late-type systems (A_{B} < 0.3), and the cosmological parameters derived from those SNe Ia hosted by (presumed) dust-free early-type galaxies are consistent with our previous determination of a non-zero Λ. The brightness scatter about the Hubble line for SNe Ia in these early-type hosts is also significantly smaller than for the SNe Ia in late-type galaxies. This result was based on HST STIS “snapshot” images and Keck spectroscopy of SNe spanning the range 0.3 < z < 0.8.

GO-8346: We had the unique opportunity of following up SN200fr, which had been discovered 14 days prior to maximum light in its restframe. Because this supernova at z=0.54 was discovered so early we were able to obtain excellent light curves from HST in F555W, F675W and F814W spanning the period from one week prior to maximum light to 6 weeks after Several spectra of the supernova were taken at VLT and Keck along with NIR photometry at VLT. To date, this is still the best observed high-redshift supernova and preliminary results were presented in Nobili, S. et al. 2001, AAS, 199,1611N.

DD-8088: WFPC2 and NICMOS (cycle 7) observations were obtained for SN1998eq at z = 1.20 (a record-breaking redshift for a spectroscopically confirmed Type Ia supernova; Aldering, et al., 1998,IAUC,7046.) The preliminary photometry is consistent with the previous results for Ω_{M}, Ω_{Λ}. With the final NICMOS image of the galaxy without the supernova recently obtained in December 2002, this analysis can now be completed.

GO-7850 and balance of GO-7336 and DD-7590: WFPC2 and NICMOS observations were obtained for 11 Type Ia supernovae in the redshift range 0.36—0.86. These observations, including final references where necessary, are now complete, and the results are about to be submitted for publication in Knop, R., et al. 2003, (in preparation). A preliminary Hubble diagram was presented January 2002 AAS meeting. The cosmological results from these SNe are in close agreement with results from the first supernova results (Per99) that gave direct evidence for a cosmological constant. The lightcurves provided by WFPC2 for these supernovae were excellent; at the higher redshifts, these lightcurves provide a substantially better measurement of the calibrated supernova magnitude than those for comparable supernovae observed only from the ground. The color information provided by NICMOS (Burns, S., et al., 2001,AAS,199.1610B), was only possible with HST. The improvement of the confidence limits on the cosmological parameters Ω_{M} and Ω_{Λ} are as good as we had previously predicted.

GO-7336 and DD-7590: Perlmutter et al., 1998, Nature, 391, 51 reported the results of our HST and ground-based imaging and Keck spectroscopic observations of SN1997ap, then the highest redshift (z = 0.83) spectroscopically confirmed Type Ia supernova. The HST portion is based on a total of 4 orbits. Also from this program, HST observations of two z = 0.83 SNe are included in the analysis in Per99 which reports on the results from our HST and ground-based imaging and Keck spectroscopic observations of 42 type Ia supernovae with 0.18 < z < 0.86. The paper rules out a flat Ω_{M} = 1 universe and presents very strong evidence for a positive cosmological constant.